OpenStudy (anonymous):

Solve D = one-fourth (e – 3f) for e. You must show all work

OpenStudy (espex):

\(D=\frac{1}{4}(e-3f)\) and you must isolate 'e'?

OpenStudy (anonymous):

can you show me step by step

OpenStudy (espex):

Certainly. Before you can do any moving on the right you must expand the expression \(\frac{1}{4}(e-3f)\), what would be the first step here?

OpenStudy (espex):

The parenthesis are telling you that \(\frac{1}{4}\) is being multiplied by both terms inside, so you want to distribute first.